A random sample of 10 chocolate energy bars of a certain brand has, on average, 230 calories per bar, with a standard deviation of 15 calories. Construct a 99% confidence interval for the true mean calorie content of this brand of energy bar. Assume that the distribution of the calorie content is approximately normal.

1
Expert's answer
2022-04-27T17:10:28-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{s}{\sqrt{n}},"


where:

  • "\bar{x}=230" : sample mean
  • zthe chosen z-value, for a 99% confidence interval z = 2.576
  • "s=15" sample standard deviation
  • n = 10 sample size.

So,

"CI=230 \pm 2.576\cdot\cfrac{15}{\sqrt{10}}=\\=
230\pm12.22=(217.78, 242.22)."